# Bayesian Deep Learning for Discrete Choice

### v1

@@ 1 Introduction
Given that inference is fundamental in the discrete choice field, researchers often rely on transparent and interpretable statistical binary or multinomial classification models such as logistic and probit regressions, along with their more complex variations. Traditional discrete choice models (DCMs) allow for point and interval estimation of key economic quantities, including marginal rates of substitution and odds ratios. However, because prediction is typically not the primary objective in discrete choice studies, these models often underperform in predictive tasks compared to machine learning approaches, particularly deep learning (DL) models [1].
@@
The first challenge, highlighted by [2], involves incorporating behavioral assumptions into model design to extract economic information comparable to standard discrete choice models. The second challenge, discussed by [3], arises from under-specification, where multiple parameter estimates yield similar predictive performance but differing economic implications. The third challenge, closely related to the second, stems from limited data availability, leading to uncertainty in model-derived hypotheses [4]. To our knowledge, the limited number of deep learning applications to discrete choice have not addressed uncertainty representation and have focused solely on point estimation. Consequently, interval estimation and hypothesis testing using deep learning models in discrete choice remain open problems.
@@
Deep learning models have been increasingly explored—albeit still to a limited extent—in discrete choice studies focused on point estimation and predictive performance. In such applications, DL models often achieve higher predictive accuracy (e.g., [1], [5], [6]) and offer greater flexibility than traditional discrete choice models (DCMs). Additionally, previous work has demonstrated that DL models can yield economic insights as comprehensive as those derived from traditional DCMs [2].
@@
Still, DL models applied to discrete choice are often highly underspecified by the data they are trained on—typically datasets with only thousands of observations—making them prone to overfitting and resulting in unstable parameter estimates, which, in turn, affects their interpretability. To address these challenges, researchers have proposed regularization techniques and inductive biases. For instance, in [7], Wang introduced a novel model design that enforces independence from irrelevant alternatives (IIA) by using shared-parameter model blocks, thereby reducing model complexity. As an example of regularization, in [8], Feng et al. proposed a gradient-based regularization technique that encourages behavioral consistency by enforcing expected gradient signs with respect to input attributes.
@@
In [9], Sifringer et al. presented a deep learning (DL) model architecture comprised of a traditional discrete choice model and a fully connected neural network. In their approach, the neural network and the traditional discrete choice model components did not share input features to ensure that the linear terms were not overrun by the nonlinearities in the neural network. Their model specification provided the first approximation of a deep learning model for discrete choice with high interpretability while achieving better predictive performance than traditional approaches. They also established a connection between multinomial logit models from discrete choice and one-dimensional convolutions from deep learning. Nevertheless, they did not address issues related to parameter instability and uncertainty representation.
@@
So far, few studies have focused on uncertainty representation, interval estimation, or hypothesis testing in deep learning for discrete choice. However, in [10], we briefly demonstrate how to construct credible intervals at the individual level using approximate Bayesian inference, attaining results that align with behavioral intuition.
@@
Arguably, when addressing the issue of uncertainty representation in deep learning, the set of tools that come to mind originates from the relatively new framework of conformal prediction. In conformal prediction, uncertainty is quantified and used to create regression bands or classification sets with strong finite-sample coverage guarantees [11]. One of its main advantages is that, in general, it requires no more than a single additional forward pass of the model through the dataset and does not depend on any model or data distribution assumptions to work as intended [12]. However, as its name implies, it provides a simple—yet effective—method for uncertainty quantification in prediction sets, but not for inference on parameters associated with the input variables.
@@
Approximate Bayesian inference methods, on the other hand, can also be used for uncertainty representation [3], both in prediction sets and for inference on parameters associated with input variables. Bayesian Deep Learning (BDL) enables the incorporation of prior beliefs, providing a principled way to integrate expert knowledge into the modeling process—an approach that would significantly enhance efforts to apply deep learning to discrete choice. Furthermore, unlike standard deep learning, BDL aims to provide a distribution over model parameters rather than point estimates [13], which can be leveraged to construct interval estimates for economic quantities such as marginal rates of substitution.
@@
Multiple methods have been developed in BDL to approximate posterior distributions, including Laplace and variational approximations, ensemble methods, and posterior sampling algorithms [13]. For instance, Maddox et al. [14] proposed Stochastic Weight Averaging Gaussian (SWAG), a method that captures uncertainty in weight space with minimal computational and memory overhead compared to stochastic gradient descent. However, it requires sampling from a high-dimensional multivariate Gaussian distribution to obtain interval prediction estimates.
@@
Other approaches based on stochastic gradient descent do not require resampling from a high-dimensional multivariate Gaussian, as in SWAG. These methods, collectively known as stochastic gradient MCMC (SG-MCMC), are algorithms that generate posterior draws directly from the training process. In this paper, we focus on stochastic gradient Langevin dynamics (SGLD) [15]. While its implementation is straightforward and widely adopted in the machine learning community, we demonstrate its effective integration and adaptation to the unique challenges of the discrete choice field. Building on this foundation, we present a simple yet general DL model architecture specifically designed to leverage the power of approximate Bayesian inference methods such as SWAG and SGLD. Our goal is to develop a DL model that collapses to behaviorally informed hypotheses when data is limited, thereby avoiding overfitting and instability in underspecified settings while remaining capable of capturing complex nonlinear relationships when sufficient data is available. We demonstrate our architecture under SGLD through a Monte Carlo simulation study, which we use to estimate predictive metrics such as out-of-sample balanced accuracy and inferential metrics such as empirical coverage for marginal rates of substitution interval estimates. In addition, we provide inference results for the stated and revealed preference data, focusing on estimates for the value of travel time savings.
@@ 2 Standard Discrete Choice Modeling
where the alternative attribute parameters $\bm{\beta}$ are shared across alternatives, while $\bm{\gamma_{j}}$ are defined only for $J-1$ alternatives to ensure identifiability. A representation of this model as a shallow neural network is provided in Figure 1. As depicted in the figure, the alternative-specific attributes undergo a 1D convolution with a kernel size equal to the number of alternative attributes, while the shared individual-specific characteristics pass through independent model blocks with parameters equal to the number of individual-specific characteristics. The relevant outputs from the 1D convolution for each alternative, along with those from the $J-1$ independent model blocks, are summed per alternative to obtain $J$ representative utility values. [^2] For alternative $J$, the representative utility depends only on the relevant output from the 1D convolution. Alternative $J$ serves as the base alternative, as all alternative parameters related to individual-specific attributes are computed relative to it, ensuring model identifiability. These representative utilities then pass through a Softmax activation function to generate probability predictions for each alternative, representing the likelihood of the individual choosing each option given their characteristics and the attributes of the alternatives.
@@
For a more detailed and general description of standard discrete choice models, refer to [16].
@@ 3 Deep Learning for Discrete Choice
The first challenge was highlighted by [2], who demonstrated how to extract economic information as comprehensive as that provided by standard discrete choice models and illustrated methods for incorporating behavioral assumptions into model specifications through architectural design.
@@
The second challenge concerns the shape of the cost function during training when the model is underspecified by the data, as discussed by [3]. Specifically, when the model architecture includes a large number of parameters but is constrained by limited data, the cost function tends to exhibit extensive valleys and multiple modes, leading to numerous local minima. This complex landscape can yield different parameter estimates with equivalent predictive performance in the training data but distinct and sometimes implausible economic interpretations.
@@
The third challenge—closely related to the second—is that, in most discrete choice inference problems, researchers work with very limited data, leading to uncertainty in the hypotheses derived from the model-fitting process [4]. To our knowledge, only a limited number of deep learning applications to discrete choice have addressed the representation of epistemic uncertainty, as most have primarily focused on point estimation.
@@
For instance, very recently, Zheng et al. [17] used Bayesian Neural Networks (BNNs) to quantify prediction uncertainty for guiding more efficient survey data collection in travel mode choice. However, they did not propose an interpretable model that could be used to derive economic insights, nor did they compute point or interval estimates for marginal utilities or other relevant information. Nevertheless, their application of BNNs demonstrates how uncertainty quantification can be leveraged in discrete choice settings, even without interpretable models.
@@
Arguably, the primary focus in the field of deep learning for discrete choice has been on designing architectures that incorporate economic assumptions and expert knowledge, leading to constrained model classes with better generalization. Some studies have proposed architectures that enforce independence from irrelevant alternatives by structuring neural networks with independent model blocks for each alternative, resulting in architectures similar–although deeper–to the one presented in Figure 1 [e.g., [2], [18], [10]].
@@
Other works have focused on augmenting discrete choice models with deep learning architectures while preserving their interpretability. The quintessential example of this approach comes from Sifringer et al. [9], who formulate a deep learning discrete choice model with two distinct components: an interpretable model informed by expert knowledge, which depends on attributes relevant for inference, and a data-driven deep learning model with input attributes that, while not relevant for inference, may still inform individual utilities.
@@
Following a similar approach, Wong et al. [19] propose a Residual Neural Network (ResLogit), in which a representative portion of the utilities depends on alternative-specific attributes, while another portion is influenced by attributes from all alternatives. This structure allows the model to learn complex substitution patterns in an interpretable manner. However, their model does not explicitly address how to incorporate attributes shared by all alternatives, such as individual characteristics. Additionally, the linear component of the utility may be overshadowed by the nonlinear component that depends on all alternative attributes.
@@
In [10], we introduce an architecture (Figures 3 and 3) that models scenarios with IIA, leverages residual connections for improved identifiability, enables automatic feature learning and attribute interactions, controls non-linearity magnitudes, incorporates individual characteristics, and captures peer effects using graph convolutional neural networks. Building on these innovations, this paper proposes a general deep learning model designed to work effectively with approximate Bayesian inference.
Figure 2: Binary Skip-GNN model architecture[10].
Figure 3: Multinomial Skip-GNN with IIA model architecture [10].
@@ 4 Approximate Bayesian Inference for Deep Learning Models
A common concern with deep learning models is their tendency to overfit training datasets due to the neglect of epistemic and aleatoric uncertainty [20]. Their flexibility allows them to represent multiple hypotheses consistent with the training data. Some studies [e.g., [21]] examine the geometry of the loss surface in deep learning models across various domains. Researchers have found that these landscapes are often highly non-convex, featuring multiple local optima and large connected valleys—where different configurations of model parameters $\bm{\Theta}$ achieve nearly the same posterior probability. However, non-Bayesian deep learning approaches rely on a single model, disregarding uncertainty that could be captured by disagreement among multiple plausible models—each with high posterior probability—resulting in poor generalization.
A key advantage of Bayesian deep learning is the use of marginalization rather than optimization [20]. The predictive distribution in a Bayesian setting is given by Equation 6, where different parameter sets $\bm{\Theta}$ are weighted by their posterior distribution $p(\bm{\Theta}|\bm{y},\bm{X},\bm{Q})$. If this integral were analytically computable, predictions would follow a Bayesian Model Average (BMA) [22].
[…]
@@
Recent applications of deep learning (DL) to discrete choice modeling suggest that model interpretability and generalization can be improved by leveraging ensemble methods [2]. Since deep learning architectures are often underspecified due to limited training data, they tend to exhibit irregular loss landscapes with large connected valleys and multiple modes [23, 24]. This results in diffuse posteriors with significant parameter uncertainty, making Bayesian approaches particularly valuable for these models. By marginalizing over a set of $\bm{\Theta}$ configurations instead of relying on a single set of model weights, Bayesian deep learning effectively captures uncertainty and improves robustness by averaging multiple high-performing models.
To address these challenges, researchers can integrate into their studies methods such as Stochastic Weight Averaging (SWA), Stochastic Weight Averaging Gaussian (SWAG) [23], and Stochastic Gradient Langevin Dynamics (SGLD) [15], which approximate the posterior while maintaining computational efficiency. Deep ensembles [25], another widely used technique, have been shown to serve as posterior approximations [20], further demonstrating the benefits of Bayesian-inspired approaches. These methods not only mitigate overfitting but also facilitate uncertainty quantification for both prediction and inference. Leveraging posterior information in this way enables the construction of credible intervals for model parameters, reinforcing the case for deep learning in discrete choice applications.

### v2

@@ 1 Introduction
Given that inference is fundamental in the discrete choice field, researchers often rely on transparent and interpretable statistical binary or multinomial classification models such as logistic and probit regressions, along with their more complex variations. Traditional discrete choice models (DCMs) allow for point and interval estimation of key economic quantities, including marginal rates of substitution and odds ratios. However, because prediction is typically not the primary objective in discrete choice studies, these models often underperform in predictive tasks compared to machine learning approaches, particularly deep learning (DL) models [wang2021comparing].
@@
The first challenge, highlighted by [wang2020deep], involves incorporating behavioral assumptions into model design to extract economic information comparable to standard discrete choice models. The second challenge, discussed by [wilson2020bayesian], arises from under-specification, where multiple parameter estimates yield similar predictive performance but differing economic implications. The third challenge, closely related to the second, stems from limited data availability, leading to uncertainty in model-derived hypotheses [wilson2020case]. To our knowledge, the limited number of deep learning applications to discrete choice have not addressed uncertainty representation and have focused solely on point estimation. Consequently, interval estimation and hypothesis testing using deep learning models in discrete choice remain open problems.
@@
Deep learning models have been increasingly explored—albeit still to a limited extent—in discrete choice studies focused on point estimation and predictive performance. In such applications, DL models often achieve higher predictive accuracy (e.g., [wang2021comparing], [han2022neural], [lu2021modeling]) and offer greater flexibility than traditional discrete choice models (DCMs). Additionally, previous work has demonstrated that DL models can yield economic insights as comprehensive as those derived from traditional DCMs [wang2020deep].
@@
Still, DL models applied to discrete choice are often highly underspecified by the data they are trained on—typically datasets with only thousands of observations—making them prone to overfitting and resulting in unstable parameter estimates, which, in turn, affects their interpretability. To address these challenges, researchers have proposed regularization techniques and inductive biases. For instance, in [wang2020deepar], Wang introduced a novel model design that enforces independence from irrelevant alternatives (IIA) by using shared-parameter model blocks, thereby reducing model complexity. As an example of regularization, in [feng2024deep], Feng et al. proposed a gradient-based regularization technique that encourages behavioral consistency by enforcing expected gradient signs with respect to input attributes.
@@
In [sifringer2018enhancing], Sifringer et al. presented a deep learning (DL) model architecture comprised of a traditional discrete choice model and a fully connected neural network. In their approach, the neural network and the traditional discrete choice model components did not share input features to ensure that the linear terms were not overrun by the nonlinearities in the neural network. Their model specification provided the first approximation of a deep learning model for discrete choice with high interpretability while achieving better predictive performance than traditional approaches. They also established a connection between multinomial logit models from discrete choice and one-dimensional convolutions from deep learning. Nevertheless, they did not address issues related to parameter instability and uncertainty representation.
@@
So far, few studies have focused on uncertainty representation, interval estimation, or hypothesis testing in deep learning for discrete choice. However, in [2025arXiv250309786V], we briefly demonstrate how to construct credible intervals at the individual level using approximate Bayesian inference, attaining results that align with behavioral intuition.
@@
Arguably, when addressing the issue of uncertainty representation in deep learning, the set of tools that come to mind originates from the relatively new framework of conformal prediction. In conformal prediction, uncertainty is quantified and used to create regression bands or classification sets with strong finite-sample coverage guarantees [angelopoulos2021gentle]. One of its main advantages is that, in general, it requires no more than a single additional forward pass of the model through the dataset and does not depend on any model or data distribution assumptions to work as intended [angelopoulos2024theoretical]. However, as its name implies, it provides a simple—yet effective—method for uncertainty quantification in prediction sets, but not for inference on parameters associated with the input variables.
@@
Approximate Bayesian inference methods, on the other hand, can also be used for uncertainty representation [wilson2020bayesian], both in prediction sets and for inference on parameters associated with input variables. Bayesian Deep Learning (BDL) enables the incorporation of prior beliefs, providing a principled way to integrate expert knowledge into the modeling process—an approach that would significantly enhance efforts to apply deep learning to discrete choice. Furthermore, unlike standard deep learning, BDL aims to provide a distribution over model parameters rather than point estimates [papamarkou2024position], which can be leveraged to construct interval estimates for economic quantities such as marginal rates of substitution.
@@
Multiple methods have been developed in BDL to approximate posterior distributions, including Laplace and variational approximations, ensemble methods, and posterior sampling algorithms [papamarkou2024position]. For instance, Maddox et al. [maddox2019simple] proposed Stochastic Weight Averaging Gaussian (SWAG), a method that captures uncertainty in weight space with minimal computational and memory overhead compared to stochastic gradient descent. However, it requires sampling from a high-dimensional multivariate Gaussian distribution to obtain interval prediction estimates.
@@
Other approaches based on stochastic gradient descent do not require resampling from a high-dimensional multivariate Gaussian, as in SWAG. These methods, collectively known as stochastic gradient MCMC (SG-MCMC), are algorithms that generate posterior draws directly from the training process. In this paper, we focus on stochastic gradient Langevin dynamics (SGLD) [welling2011bayesian]. While its implementation is straightforward and widely adopted in the machine learning community, we demonstrate its effective integration and adaptation to the unique challenges of the discrete choice field. Building on this foundation, we present a simple yet general DL model architecture specifically designed to leverage the power of approximate Bayesian inference methods such as SWAG and SGLD. Our goal is to develop a DL model that collapses to behaviorally informed hypotheses when data is limited, thereby avoiding overfitting and instability in underspecified settings while remaining capable of capturing complex nonlinear relationships when sufficient data is available. We demonstrate our architecture under SGLD through a Monte Carlo simulation study, which we use to estimate predictive metrics such as out-of-sample balanced accuracy and inferential metrics such as empirical coverage for marginal rates of substitution interval estimates. In addition, we provide inference results for the stated and revealed preference data, focusing on estimates for the value of travel time savings.
@@ 2 Standard Discrete Choice Modeling
where the alternative attribute parameters $\bm{\beta}$ are shared across alternatives, while $\bm{\gamma_{j}}$ are defined only for J1J-1 alternatives to ensure identifiability. A representation of this model as a shallow neural network is provided in Figure 1. As depicted in the figure, the alternative-specific attributes undergo a 1D convolution with a kernel size equal to the number of alternative attributes, while the shared individual-specific characteristics pass through independent model blocks with parameters equal to the number of individual-specific characteristics. The relevant outputs from the 1D convolution for each alternative, along with those from the J1J-1 independent model blocks, are summed per alternative to obtain $J$ representative utility values. [^2] For alternative $J$, the representative utility depends only on the relevant output from the 1D convolution. Alternative $J$ serves as the base alternative, as all alternative parameters related to individual-specific attributes are computed relative to it, ensuring model identifiability. These representative utilities then pass through a Softmax activation function to generate probability predictions for each alternative, representing the likelihood of the individual choosing each option given their characteristics and the attributes of the alternatives.
@@
For a more detailed and general description of standard discrete choice models, refer to [train_2009].
@@ 3 Deep Learning for Discrete Choice
The first challenge was highlighted by [wang2020deep], who demonstrated how to extract economic information as comprehensive as that provided by standard discrete choice models and illustrated methods for incorporating behavioral assumptions into model specifications through architectural design.
@@
The second challenge concerns the shape of the cost function during training when the model is underspecified by the data, as discussed by [wilson2020bayesian]. Specifically, when the model architecture includes a large number of parameters but is constrained by limited data, the cost function tends to exhibit extensive valleys and multiple modes, leading to numerous local minima. This complex landscape can yield different parameter estimates with equivalent predictive performance in the training data but distinct and sometimes implausible economic interpretations.
@@
The third challenge—closely related to the second—is that, in most discrete choice inference problems, researchers work with very limited data, leading to uncertainty in the hypotheses derived from the model-fitting process [wilson2020case]. To our knowledge, only a limited number of deep learning applications to discrete choice have addressed the representation of epistemic uncertainty, as most have primarily focused on point estimation.
@@
For instance, very recently, Zheng et al. [zheng2024incorporating] used Bayesian Neural Networks (BNNs) to quantify prediction uncertainty for guiding more efficient survey data collection in travel mode choice. However, they did not propose an interpretable model that could be used to derive economic insights, nor did they compute point or interval estimates for marginal utilities or other relevant information. Nevertheless, their application of BNNs demonstrates how uncertainty quantification can be leveraged in discrete choice settings, even without interpretable models.
@@
Arguably, the primary focus in the field of deep learning for discrete choice has been on designing architectures that incorporate economic assumptions and expert knowledge, leading to constrained model classes with better generalization. Some studies have proposed architectures that enforce independence from irrelevant alternatives by structuring neural networks with independent model blocks for each alternative, resulting in architectures similar–although deeper–to the one presented in Figure 1 [e.g., [wang2020deep], [arkoudi2023combining], [2025arXiv250309786V]].
@@
Other works have focused on augmenting discrete choice models with deep learning architectures while preserving their interpretability. The quintessential example of this approach comes from Sifringer et al. [sifringer2018enhancing], who formulate a deep learning discrete choice model with two distinct components: an interpretable model informed by expert knowledge, which depends on attributes relevant for inference, and a data-driven deep learning model with input attributes that, while not relevant for inference, may still inform individual utilities.
@@
Following a similar approach, Wong et al. [wong2021reslogit] propose a Residual Neural Network (ResLogit), in which a representative portion of the utilities depends on alternative-specific attributes, while another portion is influenced by attributes from all alternatives. This structure allows the model to learn complex substitution patterns in an interpretable manner. However, their model does not explicitly address how to incorporate attributes shared by all alternatives, such as individual characteristics. Additionally, the linear component of the utility may be overshadowed by the nonlinear component that depends on all alternative attributes.
@@
In [2025arXiv250309786V], we introduce an architecture (Figures 3 and 3) that models scenarios with IIA, leverages residual connections for improved identifiability, enables automatic feature learning and attribute interactions, controls non-linearity magnitudes, incorporates individual characteristics, and captures peer effects using graph convolutional neural networks. Building on these innovations, this paper proposes a general deep learning model designed to work effectively with approximate Bayesian inference.
Figure 2: Binary Skip-GNN model architecture[2025arXiv250309786V].
Figure 3: Multinomial Skip-GNN with IIA model architecture [2025arXiv250309786V].
@@ 4 Approximate Bayesian Inference for Deep Learning Models
A common concern with deep learning models is their tendency to overfit training datasets due to the neglect of epistemic and aleatoric uncertainty [Wilson2020]. Their flexibility allows them to represent multiple hypotheses consistent with the training data. Some studies [e.g., [Li2018]] examine the geometry of the loss surface in deep learning models across various domains. Researchers have found that these landscapes are often highly non-convex, featuring multiple local optima and large connected valleys—where different configurations of model parameters $\bm{\Theta}$ achieve nearly the same posterior probability. However, non-Bayesian deep learning approaches rely on a single model, disregarding uncertainty that could be captured by disagreement among multiple plausible models—each with high posterior probability—resulting in poor generalization.
A key advantage of Bayesian deep learning is the use of marginalization rather than optimization [Wilson2020]. The predictive distribution in a Bayesian setting is given by Equation 6, where different parameter sets $\bm{\Theta}$ are weighted by their posterior distribution $p(\bm{\Theta}|\bm{y},\bm{X},\bm{Q})$. If this integral were analytically computable, predictions would follow a Bayesian Model Average (BMA) [Raftery1997].
$p(y_{ij}|\bm{y},\bm{X},\bm{Q})=\int_{\bm{\Theta}}p(y_{ij}|\bm{X},\bm{Q},\bm{\Theta})p(\bm{\Theta}|\bm{y},\bm{X},\bm{Q})d\bm{\Theta}$ (6)
@@
Recent applications of deep learning (DL) to discrete choice modeling suggest that model interpretability and generalization can be improved by leveraging ensemble methods [wang2020deep]. Since deep learning architectures are often underspecified due to limited training data, they tend to exhibit irregular loss landscapes with large connected valleys and multiple modes [izmailov2018averaging, li2018visualizing]. This results in diffuse posteriors with significant parameter uncertainty, making Bayesian approaches particularly valuable for these models. By marginalizing over a set of $\bm{\Theta}$ configurations instead of relying on a single set of model weights, Bayesian deep learning effectively captures uncertainty and improves robustness by averaging multiple high-performing models.
To address these challenges, researchers can integrate into their studies methods such as Stochastic Weight Averaging (SWA), Stochastic Weight Averaging Gaussian (SWAG) [izmailov2018averaging], and Stochastic Gradient Langevin Dynamics (SGLD) [welling2011bayesian], which approximate the posterior while maintaining computational efficiency. Deep ensembles [Lakshminarayanan2017], another widely used technique, have been shown to serve as posterior approximations [Wilson2020], further demonstrating the benefits of Bayesian-inspired approaches. These methods not only mitigate overfitting but also facilitate uncertainty quantification for both prediction and inference. Leveraging posterior information in this way enables the construction of credible intervals for model parameters, reinforcing the case for deep learning in discrete choice applications.